# Renormalon-based resummation for spacelike and timelike QCD quantities whose perturbation expansion has general form

### v1

@@ V An example: Implementation of resummation, timelike case
It turns out that the coefficients $f_{n}(\nu_{0};\kappa)$ are known212121 In GandN the relevant coefficients are denoted as $c_{n}$, and they are related to our $f_{n}(\nu_{0};\kappa)$ ($\kappa=1$) as: $f_{n}(\nu_{0};1)=c_{n}/4^{n}$ ($n=0,1,\ldots$). Their quantity ${\hat{C}}(m_{q})$ is related to ${\cal F}(\sigma)$ as: ${\hat{C}}(m_{q})=$ $\pi^{\nu_{0}}{\cal F}(m_{q}^{2})$. for $n=0,1$ GandN :
@@
The numerical results for $f_{3}(\nu_{0};1)$ are approximate, because in order to obtain the exact value, the value of the four-loop anomalous dimension coefficient $\gamma_{3}$ is needed (which is not known). The value of $f_{3}(\nu_{0};1)$ we give here is obtained by taking $\gamma_{3}=0$. However, there are indications that this approximation is reasonably good, as it gives in the large-$\beta_{0}$ approximation a value that is close to the known large-$\beta_{0}$ value. [^22] The large-$\beta_{0}$ value is $f_{3}^{(LB)}(\nu_{0};1)\approx 56.6608\beta_{0}^{2}$, while for the higher power on $\beta_{0}$ of the exact coefficient is $f_{3}(\nu_{0};1)\approx 57.2618\beta_{0}^{2}+\mathcal{O}(\beta_{0})$ for $\gamma_{3}=0$. We stress that the large-$\beta_{0}$ approximation is recovered by taking $\gamma_{3}=\gamma_{3}^{(LB)}$.
@@
The Borel transform has then, according to GandN , [^23] The Borel $S(u)$ given in Eq. (32) of GandN corresponds, in our conventions, to our Borel via the relation: $e^{5u/3}S(u)=4\beta_{0}(d/du)B^{(\nu_{0})}[{\cal F}](u;\kappa)$ with $\kappa=1$. the following leading $u=1/2$ IR renormalon terms
@@
As indicated, the possible relative subleading corrections ${\cal O}(1/2-u)$ to this ansatz Eq. (91) will be neglected. This expression has four parameters: ${\widetilde{K}}_{q}^{(1)}$ (q=e,+,0,q=e,+,0,-).262626 The exponential factor $\exp({\widetilde{K}}_{e}^{(1)}u)$ in Eq. (91) can be interpreted as $\exp({\widetilde{K}}_{e}^{(1)}u)=\exp({\widetilde{K}}_{e}^{(1)}/2)(1-{\widetilde{K}}_{e}^{(1)}(1/2-u))$ ${\cal O}((1/2-u)^{2})$. This means that this factor generates next-to-leading relative corrections ${\cal O}(1/2-u)$ to each of the three leading renormalon terms of Eq. (89) in the way that rescaling of the momentum would generate. These four parameters can be determined by the knowledge of the original first four coefficients $f_{n}(\nu_{0};\kappa)$ (at $\kappa=1$), Eqs. (81)-(82).
@@
The resummations are performed according to the formulas (79) in pQCD and (78) in (holomorphic) AQCD, where $F_{{{\cal D}^{(1)}}}(t)_{(p,{\widetilde{s}})}$ is replaced by $F_{{{\cal D}^{(1)}}}(t)$ of Eq. (92). Specifically, the sought resummation of the considered timelike quantity ${\cal F}(\sigma)$ in AQCD is then
${\cal F}(\sigma)_{\rm res.}=\int_{0}^{\infty}\frac{dt}{t}F_{{{\cal D}^{(1)}}}(t){\widetilde{\mathfrak{H}}}_{\nu_{0}}(te^{-{\widetilde{K}}_{e}}\sigma),$ (93)
@@
Table 4: The parameters ${\widetilde{K}}_{q}$ appearing in the modified Borel transform ${\widetilde{B}}[{{\cal D}^{(1)}}](u)$ Eq. (91), for $n_{f}=3$ and $n_{f}=4$ (in the LMM scheme, and for $\kappa=1$). The corresponding scheme parameter values $c_{j}$ ($j=2,3$) and the power index $\nu_{0}$ are included.
$n_{f}$ ${\widetilde{K}}_{e}$ $\widetilde{K}$$_{+}$ ${\widetilde{K}}_{0}$ $\widetilde{K}$$_{-}$ $c_{2}$ $c_{3}$ $\nu_{0}(n_{f})$
@@
We then evaluate the timelike quantity of interest ${\cal F}(\sigma)$ with the resummation (93), for $n_{f}=3,4$. In particular, we evaluate it: (a) at $\sigma=m_{c}^{2}$ (with $n_{f}=3$) where $m_{c}=(1.67\pm 0.07)$ GeV is the pole mass of charm quark PDG2024 ; (b) at $\sigma=m_{b}^{2}$ (with $n_{f}=4$) where $m_{b}=(4.78\pm 0.06)$ GeV is the pole mass of the bottom quark PDG2024 .
@@
It turns out that this ratio is the leading order approximation for the following ratio of mass splitting between the ground-state pseudoscalar and vector mesons, in the bottom and charm quark systems GandN
@@
To improve the $\overline{\rm MS}$ TPS result in the future, it is desirable to have a parametric control on the renormalon contributions in the Hyperasymptotic approximation AP1 ; AP2 .
@@ Appendix A Proof of Theorem 2
The idea of the proof of Theorem 2 is to find the ratio between the coefficients ${\widetilde{d}}_{n}(\nu_{0};\kappa)$ and $d_{n}(\nu_{0},\kappa)$ for large $n$. For this, the relations (3) at large $n$ will be investigated here.303030 The notational conventions for the $\beta$-function coefficients $\beta_{j}$ and $c_{j}$ are given in Eqs. (6) where, as always, $a=\alpha_{s}/\pi$.
@@
As shown above around Eq. (105b), the asymptotic expression (105b) for $d_{n}(\nu_{0};\kappa)$ is equivalent to the form Eq. (25) of the Borel $B({\cal D}](u;\kappa)$. Comparing the obtained asymptotic expression (120) for ${\widetilde{d}}_{n}(\nu_{0};\kappa)^{(c_{1})}$ with that of $d_{n}(\nu_{0};\kappa)$ Eq. (105b), we see that the modified Borel ${\widetilde{B}}^{(\nu_{0})}[{\cal D}](u;\kappa)$ has the same structure, with the difference that the index $s_{0}$ is replaced by ${\widetilde{s}}_{0}=s_{0}-p\beta_{1}/\beta_{0}^{2}$.
This proves Theorem 2 for the case when the effects of the higher beta-coeffiicents are neglected (i.e., $c_{2}=c_{3}=\cdots=0$). [^31] The radiative corrections ${\cal O}(1/n)$ in the result (120) clearly correspond to $s_{0}\mapsto(s_{0}-1)$, i.e., in the modified Borel ${\widetilde{B}}^{(\nu_{0})}[{\cal D}](u;\kappa)$ they correspond to the relative corrections ${\cal O}(p-u)$, cf. Eq. (26). These corrections are in general nonzero even if the relative corrections ${\cal O}(p-u)$ in the Borel $B^{(\nu_{0})}[{\cal D}](u;\kappa)$ are taken to be zero.
@@
In the specific case considered in the main text of this work, we have $n_{f}=3$ and $p=1/2$, therefore $1/b=1/4.5=0.2222$ is small and the sum in Eq. (129) is expected to converge well.323232 For $p=1/2$ and $n_{f}=3$, the sum in the brackets on the RHS of Eq. (129) is: 10.07941-0.0794 0.0121-0.0121 $0.0005\times 0.7^{1+\beta^{(4)}}$.
@@ Appendix B Proof of Theorem 4
$\left\{\int_{-R}^{R}+\int_{{\cal C}_{+}(R)}+\int_{{\cal C}_{c}^{+}(\varepsilon)}+\int_{{\cal C}_{c}^{-}(\varepsilon)}+\int_{{\cal C}_{-}(\varepsilon)}\right\}\frac{dze^{+iz|\ln t|}}{2(iz+p)^{{\widetilde{s}}}}$. (134)

### v2

@@ V An example: Implementation of resummation, timelike case
It turns out that the coefficients $f_{n}(\nu_{0};\kappa)$ are known [^24] In GandN the relevant coefficients are denoted as $c_{n}$, and they are related to our $f_{n}(\nu_{0};\kappa)$ ($\kappa=1$) as: $f_{n}(\nu_{0};1)=c_{n}/4^{n}$ ($n=0,1,\ldots$). Their quantity ${\hat{C}}(m_{q})$ is related to ${\cal F}(\sigma)$ as: ${\hat{C}}(m_{q})=$ $\pi^{\nu_{0}}{\cal F}(m_{q}^{2})$. for $n=0,1$ GandN :
@@
The numerical results for $f_{3}(\nu_{0};1)$ are approximate, because in order to obtain the exact value, the value of the four-loop anomalous dimension coefficient $\gamma_{3}$ is needed (which is not known). The value of $f_{3}(\nu_{0};1)$ we give here is obtained by taking $\gamma_{3}=0$. However, there are indications that this approximation is reasonably good, as it gives in the large-$\beta_{0}$ approximation a value that is close to the known large-$\beta_{0}$ value. [^25] The large-$\beta_{0}$ value is $f_{3}^{\rm(LB)}(\nu_{0};1)\approx 56.6608\beta_{0}^{2}$, while the term with the highest power of $\beta_{0}$ in the exact coefficient is $f_{3}(\nu_{0};1)\approx 57.2618\beta_{0}^{2}+\mathcal{O}(\beta_{0})$ for $\gamma_{3}=0$. If we used, instead of $\gamma_{3}=0$, the values $\gamma_{3}=\gamma_{3}^{\rm(LB)}$ (=10516.,8348.=-10516.,-8348., for $n_{f}=3,4$), we would obtain $f_{3}(\nu_{0};1)=190.373,147.508$ for $n_{f}=3,4$, which are close to the case of $\gamma_{3}=0$.
@@
The Borel transform has then, according to GandN , [^26] The Borel $S(u)$ given in Eq. (32) of GandN corresponds, in our conventions, to our Borel via the relation: $e^{5u/3}S(u)=4\beta_{0}(d/du)B^{(\nu_{0})}[{\cal F}](u;\kappa)$ with $\kappa=1$. the following leading $u=1/2$ IR renormalon terms:
@@
As indicated, the possible relative subleading corrections ${\cal O}(1/2-u)$ to this ansatz Eq. (95) will be neglected. This expression has four parameters: ${\widetilde{K}}_{q}^{(1)}$ (q=e,+,0,q=e,+,0,-). [^27] The exponential factor $\exp({\widetilde{K}}_{e}^{(1)}u)$ in Eq. (95) can be interpreted as $\exp({\widetilde{K}}_{e}^{(1)}u)=\exp({\widetilde{K}}_{e}^{(1)}/2)(1-{\widetilde{K}}_{e}^{(1)}(1/2-u))$ ${\cal O}((1/2-u)^{2})$. This means that this factor generates next-to-leading relative corrections ${\cal O}(1/2-u)$ to each of the three leading renormalon terms of Eq. (93) in the way that rescaling of the momentum would generate. These four parameters can be determined by the knowledge of the original first four coefficients $f_{n}(\nu_{0};\kappa)$ (at $\kappa=1$), Eqs. (85)-(86).
@@
The resummations are performed according to the formulas (83) in pQCD and (82) in (holomorphic) AQCD, where $F_{{{\cal D}^{(1)}}}(t)_{(p,{\widetilde{s}})}$ is replaced by $F_{{{\cal D}^{(1)}}}(t)$ of Eq. (96), and ${\widetilde{K}}_{e}\mapsto{\widetilde{K}}_{e}^{(1)}$
${{\cal D}^{(1)}}(Q^{2})_{\rm res.}$ $=$ $\int_{0}^{\infty}\frac{dt}{t}F_{{{\cal D}^{(1)}}}(t){\mathcal{A}}(te^{-{\widetilde{K}}_{e}^{(1)}}Q^{2}),$ (97a)
${\cal D}(Q^{2})_{\rm res.}$ $=$ $\int_{0}^{\infty}\frac{dt}{t}F_{{{\cal D}^{(1)}}}(t){\widetilde{\mathcal{A}}}_{\nu_{0}}(te^{-{\widetilde{K}}_{e}^{(1)}}Q^{2}),$ (97b)
and specifically, the sought resummation of the considered timelike quantity ${\cal F}(\sigma)$ in AQCD is then
${\cal F}(\sigma)_{\rm res.}=\int_{0}^{\infty}\frac{dt}{t}F_{{{\cal D}^{(1)}}}(t){\widetilde{\mathfrak{H}}}_{\nu_{0}}(te^{-{\widetilde{K}}_{e}^{(1)}}\sigma),$ (98)
@@
The exchange of the order of integration, over $dt$ and $d\phi$, leading to the result (98) [cf. Eqs. (40a), (97b) and (80a)] is justified according to Fubini’s theorem. This is so because the integrand in the double integration [^28] This integrand is: $(F_{{{\cal D}^{(1)}}}(t)/t){\widetilde{\mathcal{A}}}_{\nu_{0}}(te^{-{\widetilde{K}}_{e}^{(1)}}\sigma e^{i\phi}))/(2\pi)$. is continuous in any finite rectangle $\pi\leq\phi<+\pi$ and $0\leq t\leq t_{\rm max}$, and in general we can even make the limit $t_{\rm max}\to+\infty$. In our specific case, we have $t_{\rm max}=1$, cf. Eq. (96). We also checked numerically the legitimacy of this exchange.
@@
Table 4: The parameters ${\widetilde{K}}_{q}^{(1)}$ appearing in the modified Borel transform ${\widetilde{B}}[{{\cal D}^{(1)}}](u)$ Eq. (95), for $n_{f}=3$ and $n_{f}=4$ (in the LMM scheme, and for $\kappa=1$). The corresponding scheme parameter values $c_{j}$ ($j=2,3$) and the power index $\nu_{0}$ are included.
$n_{f}$ ${\widetilde{K}}_{e}^{(1)}$ ${\widetilde{K}}_{+}^{(1)}$ ${\widetilde{K}}_{0}^{(1)}$ ${\widetilde{K}}_{-}^{(1)}$ $c_{2}$ $c_{3}$ $\nu_{0}(n_{f})$
@@
We then evaluate the timelike quantity of interest ${\cal F}(\sigma)$ with the resummation (98), for $n_{f}=3,4$. In particular, we evaluate it: (a) at $\sigma=m_{c}^{2}$ (with $n_{f}=3$) where $m_{c}=(1.67\pm 0.07)$ GeV is the pole mass of charm quark PDG2024 ; (b) at $\sigma=m_{b}^{2}$ (with $n_{f}=4$) where $m_{b}=(4.78\pm 0.06)$ GeV is the pole mass of the bottom quark PDG2024 . [^31] In this context, we note that another type of resummation for spacelike and timelike observables was constructed in Refs. SM1 ; SMnu ; SMnuaux . In SM1 it was constructed for observables with $\nu_{0}=1$: In the one-loop approximation, it turns out to give the same form as our approach in the one-loop approximation (and that of Neubert Neubert ), although it is constructed in a different way. But in the case beyond the one-loop, the resummation in SM1 has a different form which includes in the integrand, besides the mentioned one-loop-type characteristic function (i.e., the one that fulfills the sum rules (18) with $d_{n}$ instead of ${\widetilde{d}}_{n}$ on the left-hand side), the running coupling $a$ at scales which depend on that running coupling itself. The application of such resummations in general AQCD variants appears to be more complex. In Ref. SMnu (cf. also SMnuaux ) the approach of SM1 was extended to the general observables with noninteger $\nu_{0}$, in the one-loop approximation.
@@
It turns out that this ratio is the leading order approximation for the following ratio of mass splitting between the ground-state pseudoscalar and vector mesons, in the bottom and charm quark systems GandN (cf. also ABN )
@@
The quantities ${\hat{C}}(m_{q})$ can formally be considered as observables because they are renormalisation scale (RScl) and scheme (RSch) invariant. The scale $\Lambda_{\rm eff}$, appearing in the relation (108), is thus also RScl and RSch invariant, thus formally an observable.
@@
In the work GandN , the quantities ${\hat{C}}(m_{b})$ and ${\hat{C}}(m_{c})$ were calculated as well, as TPS in $\overline{\rm MS}$, with two terms in ${\hat{C}}(m_{c})$ and three terms in ${\hat{C}}(m_{b})$ (due to strong divergence). They used $\alpha_{s}^{\overline{\rm MS}}(m_{c}^{2})=0.36$ and $\alpha_{s}^{\overline{\rm MS}}(m_{b}^{2})=0.22$, and obtained for the ratio the value ${\hat{C}}(m_{b})/{\hat{C}}(m_{c})\approx 0.80$ and consequently $\Lambda_{\rm eff}\approx 0.220$ GeV. Uncertainties were not estimated there.
To improve the $\overline{\rm MS}$ TPS result in the future, it is desirable to have a parametric control on the renormalon contributions in the Hyperasymptotic approximation Ayala:2019uaw ; Ayala:2019hkn .
@@ VI Conclusions
Data availability statement:
All the necessary data that support the findings of this study are included within the article. If desired, some specific numerical programs (in Mathematica) can be obtained from one of the authors (G.C.).
@@ Appendix A Proof of Theorem 2
The idea of the proof of Theorem 2 is to find the ratio between the coefficients ${\widetilde{d}}_{n}(\nu_{0};\kappa)$ and $d_{n}(\nu_{0},\kappa)$ for large $n$. For this, the relations (3) at large $n$ will be investigated here. [^33] The notational conventions for the $\beta$-function coefficients $\beta_{j}$ and $c_{j}$ are given in Eqs. (6) where, as always, $a=\alpha_{s}/\pi$.
@@
As shown above around Eq. (110b), the asymptotic expression (110b) for $d_{n}(\nu_{0};\kappa)$ is equivalent to the form Eq. (25) of the Borel $B[{\cal D}](u;\kappa)$. Comparing the obtained asymptotic expression (125) for ${\widetilde{d}}_{n}(\nu_{0};\kappa)^{(c_{1})}$ with that of $d_{n}(\nu_{0};\kappa)$ Eq. (110b), we see that the modified Borel ${\widetilde{B}}^{(\nu_{0})}[{\cal D}](u;\kappa)$ has the same structure, with the difference that the index $s_{0}$ is replaced by ${\widetilde{s}}_{0}=s_{0}-p\beta_{1}/\beta_{0}^{2}$.
This proves Theorem 2 for the case when the effects of the higher beta-coefficients are neglected (i.e., $c_{2}=c_{3}=\cdots=0$). [^34] The radiative corrections ${\cal O}(1/n)$ in the result (125) clearly correspond to $s_{0}\mapsto(s_{0}-1)$, i.e., in the modified Borel ${\widetilde{B}}^{(\nu_{0})}[{\cal D}](u;\kappa)$ they correspond to the relative corrections ${\cal O}(p-u)$, cf. Eq. (26). These corrections are in general nonzero even if the relative corrections ${\cal O}(p-u)$ in the Borel $B^{(\nu_{0})}[{\cal D}](u;\kappa)$ are taken to be zero.
@@
In the specific case considered in the main text of this work, we have $n_{f}=3$ and $p=1/2$, therefore $1/b=1/4.5=0.2222$ is small and the sum in Eq. (134) is expected to converge well. [^35] For $p=1/2$ and $n_{f}=3$, the sum in the brackets on the RHS of Eq. (134) is: 10.07941-0.0794 0.0121-0.0121 $0.0005\times 0.7^{1+\beta^{(4)}}$.
@@ Appendix B Proof of Theorem 4
$\left\{\int_{-R}^{R}+\int_{{\cal C}_{+}(R)}+\int_{{\cal C}_{c}^{+}(\varepsilon)}+\int_{{\cal C}_{c}^{-}(\varepsilon)}+\int_{{\cal C}_{-}(\varepsilon)}\right\}\frac{dze^{+iz|\ln t|}}{2(iz+p)^{{\widetilde{s}}}}=0.$ (139)